Form the PDE of the following by eliminating arbitrary functions

phi(z^2-xy,x/z)


1
Expert's answer
2022-03-23T15:33:03-0400

"\Phi(z^2-xy,x/z)=0" (1)

We set "u=z^2-xy"; "v=\frac xz".

Then the operator d on (1) leads to:

"\displaystyle d\Phi(u,v)=\frac{\partial \Phi(u,v)}{\partial u}du+\frac{\partial \Phi(u,v)}{\partial v}dv"=0.

Thus

"0=du=d(z^2-xy)" (3)

"0=dv=d(\frac{x}{z})" (4)

From (3) and (4) we have:

"2zdz-ydx-xdy=0"

"\displaystyle \frac1zdx-\frac{x}{z^2}dz=0"

Thus getting rid of dx we obtain:

"\displaystyle dx=\frac xzdz"

"\displaystyle (2z-y\frac xz)dz=xdy"

"\displaystyle dz=\frac {dy}{\frac{2z^2-yx}{xz}}=\frac{dx}{\frac{x^2}{xz}}"

"\displaystyle \frac{dz}{xz}=\frac {dy}{2z^2-yx}=\frac{dx}{x^2}"

Finally we obtain the desired PDE for "z(x,y)":

"\displaystyle x^2\frac{\partial z}{\partial x}+(2z^2-xy)\frac{\partial z}{\partial y}=xz"

Answer: "\displaystyle x^2\frac{\partial z}{\partial x}+(2z^2-xy)\frac{\partial z}{\partial y}=xz".